OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5?

OpenStudy (anonymous):

if you plug in all that imformation into the equation \[y -y_{1} = m (x-x_{1})\] then solve for y you will have a standard y = mx +b

OpenStudy (anonymous):

Is the answer y = 5x - 3

OpenStudy (anonymous):

think you might have pluged in wrong :O did you start out with: \[y - 1 = 5(x -4)\]

OpenStudy (anonymous):

Ok i think i found my mistake

OpenStudy (anonymous):

y = 5x - 1

OpenStudy (anonymous):

i will do it step by step again :) \[y- 1 = 5(x-4)\]\[y -1 = 5x - 20\]\[y = 5x - 19\] if you see a mistake plz fix it :)

OpenStudy (anonymous):

Oh... I know where I messed up,

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

no problem :)